# On the Resilience of Multi-Agent Systems with Malicious Agents

### v1

@@ 1 Introduction
Multi-agent collaboration has further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation Liu et al. (2024a); Lee et al. (2024), math problem solving Lu et al. (2024); Liang et al. (2023b), and text translation Jiao et al. (2023); Wu et al. (2024). In such multi-agent systems, improvements are achieved by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents in a role-specific manner Chen et al. (2023b); Li et al. (2024b).
However, the decentralized nature of multi-agent systems leaves them vulnerable to faulty or malicious agents, which could undermine or destroy collaboration. Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty or malicious. In a multi-agent coding system like Camel Li et al. (2024a), a malicious coding agent could produce buggy code, causing severe errors or harmful outputs when executed by another agent.
@@
To study the impact of different structures on system resilience, we select six multi-agent collaboration systems representing three classical real-world structures: Linear Hong et al. (2023); Dong et al. (2023), Flat Li et al. (2024a); Wang et al. (2023b), and Hierarchical Chen et al. (2023b); Liang et al. (2023b). We evaluate the performance of these systems across four tasks: code generation Chen et al. (2021), math problem solving Liang et al. (2023b), translation He et al. (2020), and text evaluation Wang et al. (2023a), as shown in Fig. 1. In our results, we find that the ierarchical structure exhibits the least performance degradation at $23.6\%$, aligning with its prevalence in human societal organizational structures Mihm et al. (2010). Code generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $39.6\%$. Additionally, increasing the ratio of erroneous messages and using semantic errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. We also analyze the impact of roles and rounds on system resilience, which appears minimal.
As an initial investigation into approaches for enhancing system resilience and defending against malicious agents, we introduce two strategies, each related to a specific error introduction method. The “Inspector” agent reviews and corrects messages, mirroring AutoInject which intercepts and injects errors into messages. The “Challenger” method adds to each agent’s profile the ability to challenge received messages, mirroring AutoTransform which rewrites agents’ profiles to make them malicious. Our results demonstrate that both methods increase system resilience by a higher performance in the presence of malicious agents.
@@ A Management Science Perspective on Multi-Agent Systems
3 Methodology: Introducing Erorrs
Figure 2: Overview of our error introducing process. (a) Task information. (b) Multi-agent collaboration system without malicious agents. (c) AutoTransform modifies agent’s profile to turn it into malicious. (d) AutoInject intercepts messages between agents and adds errors into the messages.
@@ 3.1 AutoTransform: Malicious Agent Transformation
AutoTransform is an LLM-based approach that takes any agent’s profile as input and outputs a profile of a malicious agent performing the same functions but introducing stealthy errors. Drawing inspiration from how we manually convert an agent into malicious one, the design of AutoTransform follows three key steps: (1) To ensure applicability to any target agent and downstream tasks, AutoTransform first analyzes the input agent profile and extract the assigned task. This step helps to extract the task and identify potential ways to produce erroneous outputs. (2) Based on the task analysis, AutoTransform lists all possible methods to inject errors, emphasizing the need for stealth to avoid detection by other agents. (3) AutoTransform then rewrites the agent’s profile with these error-injection methods, ensuring that the original functionalities of the agent remain unchanged. An example of using AutoTransform to alter an agent’s profile is shown in Fig. 2c. The complete prompt is provided in §A.2 in the appendix.
@@ 3.2 AutoInject: Direct Error Injection
AutoInject is designed for specific tasks and agents. First, we assign the task, agent, error rates ($P_{m}$ and $P_{e}$), and error type. AutoInject then selects messages [^1] The final result message is excluded to allow system recovery. from the agent with a probability of $P_{m}$ and injects errors into $P_{e}$ of the total lines or sentences in the selected message. Errors are introduced automatically using LLMs, which receive the task introduction, error type, and the specific line or sentence. The LLMs produce erroneous lines or sentences, which replace the originals. An example of using AutoInject to modify an agent’s output into erroneous is shown in Fig. 2d. Prompts for different tasks are detailed in §A.3 the appendix.
@@ 4.1 Settings
  • Math Problem Solving: CIAR Liang et al. (2023b) presents 50 questions with hidden traps to evaluate LLMs’ Counter-Intuitive Arithmetic Reasoning abilities, requiring multi-step reasoning. Accuracy is used for evaluation.
  • Translation: CommonMT He et al. (2020) consists of paired sentences to test models’ handling of three types of commonsense reasoning, especially in ambiguous contexts. We randomly sampled 100 sentences from the most challenging type, Lexical, for our evaluation, using BLEURT-20 Sellam et al. (2020); Pu et al. (2021) for evaluation, following the practice in Liang et al. (2023b).
@@
  • Flat: Camel Li et al. (2024a) presents a framework where a “User” agent iteratively refines outputs from an “Assistant” agent, applicable across various tasks. SPP Wang et al. (2023b) uses Solo-Performance-Prompting to engage a single model into three personas for coding tasks.
  • Hierarchical: MAD Liang et al. (2023b) introduces a Multi-Agent Debate framework with two debaters and one judge to promote divergent thinking in LLMs for various tasks. AgentVerse Chen et al. (2023b) employs a dynamic recruitment process, selecting agents for multi-round collaboration as needed, utilizing four agents for our selected tasks.
Not all systems are designed to support the four tasks studied in this paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §A.1 of the Appendix. By default, we use gpt-3.5-turbo-0613 as the backbone with a temperature of zero for all the six systems, AutoTransform, AutoInject, and the defense methods in §5. We introduce one malicious agent at a time to avoid interference and facilitate essential analysis. Non-malicious agents remain unaware of the malicious agent’s presence, reflecting a realistic information-asymmetric scenario Zhou et al. (2024).
@@ 4.3 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 3b illustrates the impact of AutoTransform and AutoInject across different downstream tasks. We also present the performance of single-agent, as reported by Chen et al. (2021), Liang et al. (2023b), and Chan et al. (2023) [^2] All of the baselines use GPT-3.5, aligning with multi-agent scenarios., for a clearer comparison. The results indicate several conclusions: (1) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (2) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation, math, translation, and text evaluation. Even minor errors in the first two tasks, particularly in code generation, significantly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (3) AutoTransform and AutoInject perform similarly across most tasks, except in code generation.
@@
We now present two scenarios where deliberately introduced errors enhance system performance. (1) Double Checking: Introducing an obvious error prompts the system (i.e., other agents) to require the malicious agent to produce another message to correct the erroneous code. This process not only corrects the injected error but also fixes pre-existing errors in the original code, thereby increasing the likelihood of task completion. (2) Divergent Thinking: Systems like MAD, which incorporate a debate mechanism, may sometimes get trapped in repetitive loops due to relying on the same LLMs as their backbone, resulting in stagnant discussions. By intentionally adding significant errors that shift the original distribution, we can help agents break free from these limitations. This finding aligns with and extends the conclusions from Du et al. (2023) and Liang et al. (2023b) that agents with diverse opinions can facilitate problem solving. Additionally, this mechanism explains why AutoInject can improves performance, while AutoTransform, which lets agents produce errors themselves, cannot.
@@ 4.4 RQ3: Impact of Error Rates
Increasing the number of erroneous messages causes a larger performance drop than the number of errors within a message. Since AutoTransform lacks precise control over error rates and types, we focus on AutoInject for RQ3 and RQ4. Fig. 5a presents two experiments: one with a fixed $P_{e}=0.2$ and varying $P_{m}$ at $0.2$, $0.4$, and $0.6$, labeled “Erroneous Message;” The other with a fixed $P_{m}=0.2$ and varying $P_{e}$ at $0.2$, $0.4$, and $0.6$, labeled “Errors per Message.” The performance drop for erroneous messages is nonlinear, with the most significant decrease occurring between 00 and $0.2$. As $P_{m}$ increases, the rate of decline diminishes. Regarding the error ratio in a single message are as follows: (1) In contrast to $P_{m}$, the performance decline is nearly linear as $P_{e}$ increases. (2) As $P_{e}$ increases, performance decreases, implying that while higher error rates make errors more noticeable, the agent system struggles to correct the increasing number of errors. An exception is observed when increasing $P_{e}$ from $0.4$ to $0.6$, resulting in a performance increase in three systems (MetaGPT, Self-collab, MAD). This occurs because excessive errors in a single message become noticeable, prompting other agents to request corrections. This phenomenon highlights the importance of stealth in introducing errors.
@@ 4.6 Case Study
(b) A successful attack using distraction comments.
@@ 5 Improving System Resilience
The core idea behind our defense methods involves adding a correction mechanism within the system. We explore two variants, the “Inspector” and the “Challenger.” The “Inspector,” similar to our AutoInject, is an additional agent that intercepts all messages spread among agents, checks for errors, and corrects them. This method draws inspiration from the “Police” agent in Zhang et al. (2024).
In contrast, the “Challenger,” akin to our AutoTransform, is an additional description of functionalities added in agent profiles. This method addresses the limitation that many agents can only execute assigned tasks and may not address certain problems they encounter, although they usually have the knowledge to. By empowering agents to challenge the results of others, we enhance their problem-solving capabilities. This is because most current multi-agent systems use the same LLM as the backbone for all agents, indicating their underlying ability to partially solve tasks outside their specialization. Detailed prompts for the “Police” agent and “Challenger” method can be found in §A.4 and §A.5 in the appendix.
Figure 8: Performance of defense methods, “Inspector” and “Challenger,” on code generation task.
@@ Results
We apply these defense strategies to the two weaker architectures: the linear (Self-collab) and the flat (Camel). Results for the vanilla model, error injection with AutoInject, and the two defense methods are shown in Fig. 8. Both defense methods improve performance against AutoInject, though they do not restore it to the original level. Additionally, there is no definitive conclusion as to which method is superior. We recommend trying both in practice.
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play Wang et al. (2024). Despite utilizing a same architecture, like GPT-3.5, distinct tasks benefit from tailored in-context role-playing prompts Min et al. (2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities. ChatEval Chan et al. (2023) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev Qian et al. (2023) uses a linear structure of several roles to address code generation tasks. AutoGen Wu et al. (2023) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents Chen et al. (2023a) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Agents Zhou et al. (2023a) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. There is also work simulating daily life and conversations Park et al. (2023); Zhou et al. (2023b). Additionally, there are studies on multi-agent competition Huang et al. (2024); Liu et al. (2024b); Liang et al. (2023a). These frameworks are not selected either because they are not task-oriented (e.g.., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
PsySafe Zhang et al. (2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) Tian et al. (2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2021), TruthfulQA Lin et al. (2022), MedMCQA Pal et al. (2022), and LegalBench Guha et al. (2024), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, Amayuelas et al. (2024) and Ju et al. (2024) do not explore how different system architectures are affected by these adversaries.

### v2

@@ 1 Introduction
Multi-agent collaboration has further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation Liu et al. (2023); Lee et al. (2024), math problem solving Lu et al. (2024); Liang et al. (2024), and text translation Jiao et al. (2023); Wu et al. (2024). In such multi-agent systems, improvements are achieved by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents in a role-specific manner Chen et al. (2024); Li et al. (2024).
However, the decentralized nature of multi-agent systems leaves them vulnerable to faulty or malicious agents, which could undermine or destroy collaboration. Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty or malicious. In a multi-agent coding system like Camel Li et al. (2023), a malicious coding agent could produce buggy code, causing severe errors or harmful outputs when executed by another agent.
@@
To study the impact of different structures on system resilience, we select six multi-agent collaboration systems representing three classical real-world structures: Linear Hong et al. (2024); Dong et al. (2024), Flat Li et al. (2023); Wang et al. (2024c), and Hierarchical Chen et al. (2024); Liang et al. (2024). We evaluate the performance of these systems across four tasks: code generation Chen et al. (2021), math problem solving Liang et al. (2024), translation He et al. (2020), and text evaluation Wang et al. (2024a), as shown in Fig. 1. In our results, we find that the ierarchical structure exhibits the least performance degradation at $23.6\%$, aligning with its prevalence in human societal organizational structures Mihm et al. (2010). Code generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $39.6\%$. Additionally, increasing the ratio of erroneous messages and using semantic errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. We also analyze the impact of roles and rounds on system resilience, which appears minimal.
As an initial investigation into approaches for enhancing system resilience and defending against malicious agents, we introduce two strategies, each related to a specific error introduction method. The “Challenger” method adds to each agent’s profile the ability to challenge received messages, mirroring AutoTransform which rewrites agents’ profiles to make them malicious. The “Inspector” agent reviews and corrects messages, mirroring AutoInject which intercepts and injects errors into messages. We apply these defense methods to the two weaker systems: Self-collab with a linear structure and Camel with a flat structure. Our results demonstrate that both methods enhance system resilience, recovering up to $87.9\%$ of performance lost due to malicious agents.
@@ A Management Science Perspective on Multi-Agent Systems
3 Methodology: Introducing Errors
Figure 2: Overview of our error introducing process. (a) Task information. (b) Multi-agent collaboration system without malicious agents. (c) AutoTransform modifies agent’s profile to turn it into malicious with preserving original functionalities. (d) AutoInject intercepts messages between agents and adds errors into the messages.
@@ 3.1 AutoTransform: Malicious Agent Transformation
AutoTransform is an LLM-based approach that takes any agent’s profile as input and outputs a profile of a malicious agent performing the same functions but introducing stealthy errors. Drawing inspiration from how we manually convert an agent into malicious one, the design of AutoTransform follows three key steps: (1) To ensure applicability to any target agent and downstream tasks, AutoTransform first analyzes the input agent profile and extract the assigned task. This step helps to extract the task and identify potential ways to produce erroneous outputs. (2) Based on the task analysis, AutoTransform lists all possible methods to inject errors, emphasizing the need for stealth to avoid detection by other agents. (3) AutoTransform then rewrites the agent’s profile with these error-injection methods, ensuring that the original functionalities of the agent remain unchanged. An example of using AutoTransform to alter an agent’s profile is shown in Fig. 2c. The complete prompt is provided in §B.3 in the appendix.
@@ 3.2 AutoInject: Direct Error Injection
AutoInject is designed for specific tasks and agents. First, we assign the task, agent, error rates ($P_{m}$ and $P_{e}$), and error type. AutoInject then selects messages [^1] The final result message is excluded to allow system recovery. from the agent with a probability of $P_{m}$ and injects errors into $P_{e}$ of the total lines or sentences in the selected message. Errors are introduced automatically using LLMs, which receive the task introduction, error type, and the specific line or sentence. The LLMs produce erroneous lines or sentences, which replace the originals. An example of using AutoInject to modify an agent’s output into erroneous is shown in Fig. 2d. Prompts for different tasks are detailed in §B.4 the appendix.
@@ 4.1 Settings
  • Math Problem Solving: CIAR Liang et al. (2024) presents 50 questions with hidden traps to evaluate LLMs’ Counter-Intuitive Arithmetic Reasoning abilities, requiring multi-step reasoning. Accuracy is used for evaluation.
  • Translation: CommonMT He et al. (2020) consists of paired sentences to test models’ handling of three types of commonsense reasoning, especially in ambiguous contexts. We randomly sampled 100 sentences from the most challenging type, Lexical, for our evaluation, using BLEURT-20 Sellam et al. (2020); Pu et al. (2021) for evaluation, following the practice in Liang et al. (2024).
@@
  • Flat: Camel Li et al. (2023) presents a framework where a “User” agent iteratively refines outputs from an “Assistant” agent, applicable across various tasks. SPP Wang et al. (2024c) uses Solo-Performance-Prompting to engage a single model into three personas for coding tasks.
  • Hierarchical: MAD Liang et al. (2024) introduces a Multi-Agent Debate framework with two debaters and one judge to promote divergent thinking in LLMs for various tasks. AgentVerse Chen et al. (2024) employs a dynamic recruitment process, selecting agents for multi-round collaboration as needed, utilizing four agents for our selected tasks.
Not all systems are designed to support the four tasks studied in this paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §B.1 of the Appendix. We use GPT-3.5 and GPT-4o as the backbone with a temperature of zero for main experiments (RQ1 and RQ2) while using GPT-3.5 for the remaining. Our findings are consistent with the GPT-4o results; see §A.1 in the appendix for details, as they are omitted here due to space constraints. We introduce one malicious agent at a time to avoid interference and facilitate essential analysis. Non-malicious agents remain unaware of the malicious agent’s presence, reflecting a realistic information-asymmetric scenario Zhou et al. (2024a).
@@ 4.3 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 3b illustrates the impact of AutoTransform and AutoInject across different downstream tasks. We also present the performance of single-agent using GPT-3.5 with the prompts listed in §B.2, for a clearer comparison. The results indicate several conclusions: (1) Multi-agent systems can outperform single-agent settings, but their performance may decline to similar or worse levels when affected by malicious agents. (2) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (3) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation, math, translation, and text evaluation. Even minor errors in the first two tasks, particularly in code generation, significantly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (4) AutoTransform and AutoInject perform similarly across most tasks, except in code generation.
@@
We now present two scenarios where deliberately introduced errors enhance system performance. (1) Double Checking: Introducing an obvious error prompts the system (i.e., other agents) to require the malicious agent to produce another message to correct the erroneous code. This process not only corrects the injected error but also fixes pre-existing errors in the original code, thereby increasing the likelihood of task completion. (2) Divergent Thinking: Systems like MAD, which incorporate a debate mechanism, may sometimes get trapped in repetitive loops due to relying on the same LLMs as their backbone, resulting in stagnant discussions. By intentionally adding significant errors that shift the original distribution, we can help agents break free from these limitations. This finding aligns with and extends the conclusions from Du et al. (2024) and Liang et al. (2024) that agents with diverse opinions can facilitate problem solving. Additionally, this mechanism explains why AutoInject can improves performance, while AutoTransform, which lets agents produce errors themselves, cannot.
@@ 4.4 RQ3: Impact of Error Rates
Increasing the number of erroneous messages causes a larger performance drop than the number of errors within a message. Since AutoTransform lacks precise control over error rates and types, we focus on AutoInject for RQ3 and RQ4. Fig. 5a presents two experiments: one with a fixed $P_{e}=0.2$ and varying $P_{m}$ at $0.2$, $0.4$, and $0.6$, labeled “Erroneous Message;” The other with a fixed $P_{m}=0.2$ and varying $P_{e}$ at $0.2$, $0.4$, and $0.6$, labeled “Errors per Message.” For “Erroneous Message,” as $P_{m}$ increases, the task performance consistently decreases. Regarding the error ratio in a single message: (1) In contrast to $P_{m}$, the performance reached a bottleneck as $P_{e}$ increases from $0.4$ to $0.6$. (2) As $P_{e}$ increases, performance decreases, implying that while higher error rates make errors more noticeable, the agent system struggles to correct the increasing number of errors. An exception is observed when increasing $P_{e}$ from $0.4$ to $0.6$, resulting in a performance increase in three systems (MetaGPT, Self-collab, MAD). This occurs because excessive errors in a single message become noticeable, prompting other agents to request corrections. This phenomenon highlights the importance of stealth in introducing errors.
@@ 4.6 Case Study
(b) A successful attack w/ distraction comments.
@@ 5 Improving System Resilience
The core idea behind our defense methods involves adding a correction mechanism within the system. We explore two variants, the “Challenger” and the “Inspector.” The “Challenger,” akin to our AutoTransform, is an additional description of functionalities added in agent profiles. This method addresses the limitation that many agents can only execute assigned tasks and may not address certain problems they encounter, although they usually have the knowledge to. By empowering agents to challenge the results of others, we enhance their problem-solving capabilities. This is because most current multi-agent systems use the same LLM as the backbone for all agents, indicating their underlying ability to partially solve tasks outside their specialization.
In contrast, the “Inspector,” similar to our AutoInject, is an additional agent that intercepts all messages spread among agents, checks for errors, and corrects them. This method draws inspiration from the “Police” agent in Zhang et al. (2024). Detailed prompts for the “Challenger” and “Inspector” methods can be found in §B.5 and §B.6, respectively, in the appendix.
Figure 8: Performance of defense methods, “Challenger” and “Inspector,” on code generation task.
@@ Results
We apply these defense strategies to the two weaker architectures: the linear (Self-collab) and the flat (Camel). Results for the vanilla model, error injection with AutoInject, and the two defense methods are shown in Fig. 8. Both defense methods improve performance against AutoInject, though they do not restore it to the original level. With Challenger, we recover $87.93\%$ of the performance loss caused by malicious agents (improving from $40.85\%$ to $71.95\%$, compared to the original $76.22\%$). However, no definitive conclusion can be drawn regarding which method is superior, as Inspector outperforms Challenger on Camel. We recommend testing both methods in practice.
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play Wang et al. (2024b). Despite utilizing a same architecture, like GPT-3.5, distinct tasks benefit from tailored in-context role-playing prompts Min et al. (2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities. ChatEval Chan et al. (2024) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev Qian et al. (2024) uses a linear structure of several roles to address code generation tasks. AutoGen Wu et al. (2023) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents Chen et al. (2023) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Zhou et al. (2023) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. Additionally, there are studies simulating daily life or conversations Park et al. (2023); Zhou et al. (2024b), and multi-agent competition Huang et al. (2024); Liu et al. (2024); Liang et al. (2023). These frameworks are not selected either because they are not task-oriented (e.g., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
PsySafe Zhang et al. (2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) Tian et al. (2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Agent Smith Gu et al. (2024) showed that malicious behaviors can spread among agents, using multi-agent interaction and memory storage. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2021), TruthfulQA Lin et al. (2022), MedMCQA Pal et al. (2022), and LegalBench Guha et al. (2023), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, Amayuelas et al. (2024) and Ju et al. (2024) do not explore how different system architectures are affected by these adversaries.